tom took 10 hours. together tom and jerry took 6 hours. how long would it take for jerry to finish by himself
Answer by CubeyThePenguin(3113)

Tom finishes 1/10 of the task in 1 hour.
Jerry takes 1/x of the task in 1 hour, so he can finish the task in x hours.

1/10 + 1/x = 1/6
6/60 + 1/x = 10/60
1/x = 1/15

x = 15, so Jerry can finish the task himself in 15 days.